X is the number of dollars gained or lost each day playing cards. The distribution of X is below. Give the mean and variance of X. (Show work)

X P(X)

0 .30

1 .10

2 .50

3 .10
Answer #1

Solution:

x P(x) x * P(x) x2 * P(x)
0 0.3 0 0
1 0.1 0.1 0.1
2 0.5 1 2
3 0.1 0.3 0.9
Sum 1 1.4 3

Mean = X * P(X)

= 1.4

Variance

= X 2 * P(X) - 2

= 3 - (1.4)2 = 1.04